If the nominal interest rate per year is 16,5% per annum, ompounded at the end of every se ond month, then the ee tive interest rate is

[1] 18,00%.

[2] 16,18%.

[3] 14,53%.

[4] 16,68%.

[5] 17,68%


1
Expert's answer
2021-06-13T15:45:29-0400

i=nominal interest rate

n=compounding months

r=effective interest rate

"r=(1+i/n)^n-1=(1+0.165/6)^6-1=0.1768=17.68\ \%"


Answer: 5. "17.68\ \%"